OpenStudy (anonymous):

the area of a rectangle is 9x^2-9x-4 cm^2. find its length and width then solve the perimeter Am I missing something here? The roots are -1/3 and 4/3 but I can't seem to figure out how they would come out as the length & width PLEASE HELP

OpenStudy (alekos):

from your roots that would factor out to (3x+1)(3x-4) so what do you think the length and width would be

OpenStudy (anonymous):

-1/3 & 4/3? But one of them is negative? :(

OpenStudy (anonymous):

uhm ok so the length should be (3x+1) width is (3x-1)? so i'll just substitute the 4/3 here?

OpenStudy (anonymous):

@alekos :) thanks!

OpenStudy (alekos):

no width is 3x-4

OpenStudy (alekos):

for perimeter just add l + w

OpenStudy (anonymous):

but if x = -1/3 and 4/3 3(4/3) + 1 = 5 3(4/3) - 4 = 0 ???